# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Full Static Quantization

### v1

@@ Abstract
Multimodal large language models (MLLMs) have garnered widespread attention due to their ability to understand multimodal input. However, their large parameter sizes and substantial computational demands severely hinder their practical deployment and application. While quantization is an effective way to reduce model size and inference latency, its application to MLLMs remains underexplored. In this paper, we propose MQuant, a post-training quantization (PTQ) framework designed to tackle the unique challenges of multimodal large language models (MLLMs). Conventional quantization often struggles with MLLMs because of (a) high inference latency from large visual token counts, (b) distributional disparities between visual and textual tokens, and (c) extreme outliers introduced by Hadamard-based transformations. To address these issues, MQuant introduces: • Modality-Specific Static Quantization (MSQ), assigning distinct static scales for visual vs. textual tokens; • Attention-Invariant Flexible Switching (AIFS), reordering tokens to preserve casual attention while eliminating expensive token-wise scale computations; • Rotation Magnitude Suppression (RMS), mitigating weight outliers arising from online Hadamard rotations. On five mainstream MLLMs (including Qwen-VL, MiniCPM-V, CogVLM2), MQuant under W4A8 achieves near-floating-point accuracy (<1% degradation) while reducing inference latency by up to 30%, significantly outperforming existing PTQ baselines. Our MQuant effectively bridges the gap for efficient and accurate MLLMs inference in resource-constrained devices. code will be released.
1 Introduction
Recent advances in large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b; Dubey et al., 2024) have led to remarkable performance on a wide range of natural language processing tasks. However, these models often struggle when dealing with non-textual data such as images or videos. Multimodal large language models (MLLMs) (Reid et al., 2024; Achiam et al., 2023b; Wang et al., 2023) address this limitation by integrating visual and textual modalities, thereby enabling more comprehensive understanding and reasoning. Despite these benefits, their large parameter sizes, coupled with substantial computational demands, poses a major challenge for real-world deployment, particularly in resource-constrained or privacy-sensitive environments.
@@ 1 Introduction
Figure 1: (a) The number of prefill visual tokens across different MLLMs as the image splits or resolution increases. (b) The activation values of visual tokens range from $-20$ to $10$, whereas textual tokens are centered around 00, with only a few having absolute magnitudes exceeding $0.5$.
Challenges in MLLM Quantization. Quantization has proven an effective strategy for reducing memory usage and inference costs in LLMs (Yuan et al., 2023a; Shao et al., 2023), by converting high-precision parameters (e.g., FP32) into lower-bit representations (e.g., INT8). Yet, the transition from LLMs to MLLMs brings three unique difficulties:
  1. 1.
    Time-to-First-Token (TTFT) Explosion. MLLMs often generate large numbers of visual tokens (e.g., patch embeddings or region proposals) with the resolution and aspect ratio of input images or videos. As shown in Fig 1(a), in models such as Qwen2-VL (Wang et al., 2024), the number of prefill visual tokens grows as image resolution increases (detailed figure in Fig. 7). As image or video resolution increases, the prefill visual tokens can escalate dramatically—leading to high TTFT and negatively impacting latency-sensitive tasks. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating the increase in TTFT. Per-token dynamic quantization, though flexible, exacerbates this overhead by computing scale factors token-by-token.
  2. 2.
    Disparate Modality Distributions. As shown in Fig 1 (b), the activation distributions between visual and textual tokens reveal substantial numerical discrepancies, where visual token activations can span a significantly broader range (e.g., $-20$ to $10$) than textual tokens, which typically center near 00. A single global scale factor for both visual and textual tokens leads to either aggressive clipping of visual outliers or increased quantization granularity for text, harming overall accuracy.
  3. 3.
    Visual Outliers. High-magnitude outliers in visual tokens, often reflecting salient image regions, make traditional clipping-based methods unsuitable. Naively clipping these outliers can cause severe accuracy drops, illustrated in Table 1.
Table 1: Ablations on the clip range for different tokens.
@@
Our Approach. To tackle these challenges, we propose MQuant, an accurate and efficient post-training quantization (PTQ) framework specifically tailored to MLLMs. Our contributions center around:
  1. 1.
    Modality-Specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS), which apply distinct per-tensor scales for visual vs. textual tokens and reorder multimodal inputs to avoid repeated, token-wise scale computation. By doing so, MQuant slashes TTFT and preserves accuracy.
  2. 2.
    Rotation Magnitude Suppression (RMS) for mitigating fresh outliers introduced by online Hadamard rotations. Our theoretical analysis reveals the emergence of large-mean channels in MLLM weights, and RMS effectively reduces these outliers with minimal overhead.
Experimental Results and Impact. We evaluate MQuant on five mainstream MLLMs—InternVL (Chen et al., 2024a), Qwen-VL (Bai et al., 2023b), MiniCPM-V (Yao et al., 2024), CogVLM2 (Hong et al., 2024), and Qwen2-VL (Wang et al., 2024)—over multiple vision-language benchmarks. Under a W4A8 setting, MQuant routinely maintains $>98\%$ of the floating-point accuracy, significantly outperforming baseline quantization methods. In addition, our modality-specific approach and token reordering reduce inference latency by effectively handling thousands of visual tokens at scale. We will release code to facilitate further development and encourage real-world adoption in resource-limited scenarios.
In summary, MQuant is, to our knowledge, the first PTQ framework explicitly designed for MLLMs, demonstrating both near-lossless performance and reduced computational demands at lower bit-widths. We hope our method paves the way for broader, faster, and more efficient multimodal LLM deployments.
2 Preliminaries
2.1 MLLMs Architecture and Models
As shown in the Figure, the existing MLLMs framework  (Bai et al., 2023b; Chen et al., 2024b) mainly consists of three modules: a vision encoder E for processing visual inputs, a visual-language projector P that serves as a bridge to align the two modalities and a large language model ($LLM$) that handles the multi-modal tokens and performs reasoning.
@@ 2.1 MLLMs Architecture and Models
where $l$ denotes the length of $\bm{O}_{a}$. The parameter sizes of large language models (LLMs) range from 3 billion to tens of billions. Commonly used open-source LLMs include the Llama series (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a), InternLM (Cai et al., 2024), MiniCPM (Hu et al., 2024a), ChatGLM (GLM et al., 2024).
@@
These foundational elements of MLLMs have benefited significantly from the rapid advancements in LLM technology. Flamingo (Alayrac et al., 2022) pioneered connecting pre-trained visual encoders to LLMs, demonstrating strong generalization across visual-language tasks. Following the emergence of ChatGPT (Achiam et al., 2023a), numerous open-source models based on pre-trained LLMs like LLaMA (Touvron et al., 2023a) and its variants (Touvron et al., 2023b; Zheng et al., 2024) have been proposed (Li et al., 2023; Huang et al., 2024; Zhu et al., 2023; Liu et al., 2024). Subsequent efforts like Qwen-VL (Bai et al., 2023b), InternVL (Chen et al., 2024b), and CogVLMV2 (Hong et al., 2024) enhanced MLLMs from perspectives such as high-resolution input and larger-scale training data. However, the substantial parameters of MLLMs lead to high computational costs, limiting broader application. Recently, smaller MLLMs like Mini-Gemini (Li et al., 2024), MobileVLM (Chu et al., 2024), and MiniCPM-V (Yao et al., 2024) have emerged. Despite these developments, dedicated quantization methods for MLLMs to further reduce memory usage and accelerate inference remain under-explored.
2.2 Post Training Quantization (PTQ)
@@
PTQ serves as a potent strategy for model compression. By converting the high-precision variables of pre-trained models into low-bit integers, it achieves a reduction in memory usage and an acceleration of inference speed. For uniform quantization, given a floating-point (FP) tensor $x$ (weights or activations), it can be uniformly quantized to $b$-bits in signed quantization as follows:
$\hat{\mathbf{x}}=\mathbf{Q}_{U}(\mathbf{x},b)=(clamp(\lfloor\frac{\mathbf{x}}{s}\rceil+z,q_{min},q_{max})-z)\cdot s\vspace{-2mm}$ (2)
@@
Existing post-training quantization (PTQ) methods for LLMs are categorized into weight-only and weight-activation quantization (Zhao et al., 2023; Yuan et al., 2024; Shang et al., 2023; Yue et al., 2024; Hu et al., 2024b, 2025). Weight-only methods like GPTQ (Frantar et al., 2022), QuIP (Chee et al., 2024), and AWQ (Lin et al., 2023) achieve high compression rates but offer limited inference acceleration. In contrast, weight-activation quantization methods (Xiao et al., 2022; Wei et al., 2022; Yuan et al., 2023b; Zhang et al., 2024) quantize both weights and activations, improving memory usage and latency. The main challenge is activation outliers causing quantization errors. Techniques like SmoothQuant (Xiao et al., 2022) shift quantization difficulty from activations to weights, while OmniQuant (Shao et al., 2023) optimizes performance by training quantization parameters. SliceGPT (Ashkboos et al., 2024a) reduces memory demands by designing a Pre-LN + Rotate Scheme for LLMs sparsification based on computational invariance. They achieve this by adding a linear layer in the residual connection (see Appendix  A.14). Unlike SliceGPT, we further develop a Post-LN + Rotate scheme to accommodate more vision encoder and extends its applicability to various MLLMs. This enhancement broadens the the LayerNorm + Rotate approach, making it suitable for both Pre-LN and Post-LN configurations across various MLLMs. Recently, Quarot (Ashkboos et al., 2024b) introduces rotations to eliminate outliers; however, this solution is not applicable to MLLMs due to inherent modality differences.
3 Method
In this section, we present MQuant, a post-training quantization solution specifically designed for MLLMs. In Sec. 3.1, we describe modality-specific static quantization (MSQ) and attention-invariant flexible switching (AIFS). In Sec. 3.2, we identify the weight outliers caused by the online Hadamard rotations and state Rotation Magnitude Suppression (RMS). We provide the detailed MQuant algorithm for FP MLLMs in Appendix A.1 Algorithm 1.
3.1 Modality-Specific Static Quantization and Attention-Invariant Flexible Switching
Motivation and Overview. Multi-modal Large Language Models (MLLMs) often generate a large number of visual tokens. As discussed in Section 1, the number of these tokens increases rapidly with higher image or video resolution (see Fig. 1(a)). This growth causes high inference latency when using per-token dynamic quantization, due to expensive token-wise scale computations. Such computations are not ideal for mobile or embedded devices. In contrast, static per-tensor quantization bypasses the need for repeated scale updates and offers much lower overhead. However, naïvely applying per-tensor quantization to both text and visual tokens can degrade accuracy, because their activation distributions differ significantly (Fig. 1(b)).
Figure 2: Overview of Modality-Specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS). AIFS reorders tokens so that all visual tokens appear first, then the textual tokens, while adjusting the causal mask to preserve the original model logic.
We address this issue with a two-part method: (1) Modality-Specific Static Quantization (MSQ) and (2) Attention-Invariant Flexible Switching (AIFS). MSQ applies different static scaling factors for visual and textual tokens, while AIFS reorders their positions to avoid irregular data slicing. Our method retains accuracy while benefiting from the efficiency of per-tensor quantization, especially for high-resolution inputs.
Modality-Specific Static Quantization (MSQ).
$E\;=\;\underbrace{(e^{t}_{1},\dots,e^{t}_{m-1})}_{\textstyle\text{text scale }s_{t}}\;\;\underbrace{(e^{v}_{m},\dots,e^{v}_{n})}_{\textstyle\text{visual scale }s_{v}}\;\;\underbrace{(e^{t}_{n+1},\dots,e^{t}_{L})}_{\textstyle\text{text scale }s_{t}}.$ (3)
Here, $s_{v}$ denotes the scale factor for visual tokens, while $s_{t}$ is for textual tokens. By calibrating $s_{v}$ and $s_{t}$ once before inference, we avoid the overhead of per-token scale estimation. This design aligns well with hardware that favors simple per-tensor quantization. It also prevents large visual values from saturating the narrower distribution of textual tokens.
Attention-Invariant Flexible Switching (AIFS).
Despite the benefits of MSQ, it complicates data handling if textual and visual tokens remain interleaved in $E$. Naïve slicing and concatenating at runtime can increase memory traffic and reduce efficiency for GEMM-based layers like QK and FC. To overcome this, we propose Attention-Invariant Flexible Switching (AIFS). Figure 2 illustrates the idea: we reorder the sequence so that all visual tokens appear first, followed by textual tokens. We then alter the causal mask to preserve the original auto-regressive relationships.
Causal attention (Vaswani et al., 2017) ensures each token can only attend to earlier tokens. Formally, for a naive sequence $E$, the attention matrix is
$\mathbf{A}=\mathrm{Softmax}\!\bigl{(}\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{D}}+M_{i,j}\bigr{)},\quad M_{i,j}=\begin{cases}0,&\text{if }j\leq i,\\ -\infty,&\text{otherwise}.\end{cases}$ (4)
@@ 3.1 Modality-Specific Static Quantization and Attention-Invariant Flexible Switching
With AIFS, we reorder $E$ into $E^{u}=\{\;e^{v}_{m},\dots,e^{v}_{n},\;e^{t}_{1},\dots,e^{t}_{m-1},\,\dots,e^{t}_{n+1},\dots,e^{t}_{L}\}$ and obtain a unified causal mask $M^{u}_{i,j}$:
$M^{u}_{i,j}=\begin{cases}0&\text{if one of the following conditions is met:}\\ &(i\leq(n-m),j\leq i\text{ or }(n-m)<j\leq n)\\ &\text{or }((n-m)<i\leq n,(n-m)<j\leq i)\\ &\text{or }(i>n,j\leq i)\\ -\infty&\text{otherwise}\end{cases}$ (5)
The position embeddings are also shifted consistently. More details can be found in the Appendix A.2. This guarantees that attention scores remain numerically equivalent to the original sequence. Thus, token reordering does not change the auto-regressive logic, but it does simplify memory operations so that we can feed all visual tokens through a single static-scale multiplication and do the same for textual tokens.
Efficiency and Practical Benefits. Our experiments in Section 4 show that MSQ plus AIFS delivers three main advantages:
  • High Compatibility and Strong Theoretical Equivalence: The causal-mask transformation ensures that the model’s output is numerically the same as if tokens were not reordered. This reordering can be integrated into existing MLLM implementations without major code changes.
  • Reduced Inference Latency: We replace per-token dynamic quantization with a static per-modality approach. This removes the need to recalculate scales per token, cutting runtime overhead and boosting throughput.
  • Enhanced Memory Efficiency: Placing all visual tokens first avoids repeated slicing or padding of interleaved tokens. We can handle them with a single global scale factor ($s_{v}$), then switch to $s_{t}$ for textual tokens. This reduces memory usage. We observe up to 24.7% speedups and 152.9% memory gains (see Table 11).
@@
In summary, MSQ + AIFS offers a straightforward and effective way to handle the unique challenges of multi-modal token distributions. It is especially valuable on edge devices or other constrained platforms where dynamic quantization is hard to implement at scale, yet per-tensor static quantization is hardware-friendly and fast.
3.2 FHT-Based Weight Outlier Mitigation: Rotation Magnitude Suppression
Background and Motivation.
Chee et al al. (Chee et al., 2024) and Tseng et al. (Tseng et al., 2024) introduced incoherence to measure the difficulty of quantization. A lower incoherence indicates a simpler quantization scenario. Concretely, let $W\in\mathbb{R}^{m\times n}$ be a weight matrix with singular vectors $e_{i}$ and $e_{j}$. They define
$|W_{ij}|\;=\;\bigl{|}\,e_{i}^{\top}W\,e_{j}\bigr{|}\;\;\leq\;\;\mu\,\frac{\|W\|_{F}}{\sqrt{mn}},$ (6)
where $\mu$ is the incoherence coefficient. The smaller $\mu$ is, the easier it is to quantize $W$. They also showed that applying a Hadamard transform to both weights and activations can effectively reduce $\mu$. Quarot (Ashkboos et al., 2024b) further applies offline Hadamard transforms and a partially online Hadamard transform, as shown in Figure 3(a), to achieve state-of-the-art quantization on LLMs.

### v2

# MQuant: Unleashing the Inference Potential of Multimodal Large Language Models via Static Quantization

@@ Abstract.
Multimodal large language models (MLLMs) have garnered widespread attention due to their ability to understand multimodal input. However, their large parameter sizes and substantial computational demands severely hinder their practical deployment and application. While quantization is an effective way to reduce model size and inference latency, its application to MLLMs remains underexplored. In this paper, we propose MQuant, a post-training quantization (PTQ) framework designed to tackle the unique challenges of multimodal large language models (MLLMs). Conventional quantization often struggles with MLLMs because of (a) high inference latency from large visual token counts, (b) distributional disparities between visual and textual tokens, and (c) extreme outliers introduced by Hadamard-based transformations. To address these issues, MQuant introduces: • Modality-Specific Static Quantization (MSQ), assigning distinct static scales for visual vs. textual tokens; • Attention-Invariant Flexible Switching (AIFS), reordering tokens to preserve casual attention while eliminating expensive token-wise scale computations; • Rotation Magnitude Suppression (RMS), mitigating weight outliers arising from online Hadamard rotations. On five mainstream MLLMs (including Qwen-VL, MiniCPM-V, CogVLM2), MQuant under W4A8 achieves near-floating-point accuracy (<1% degradation) while reducing inference latency by up to 30%, significantly outperforming existing PTQ baselines. Our MQuant effectively bridges the gap for efficient and accurate MLLMs inference in resource-constrained devices. Code is released in https://github.com/StiphyJay/MQuant.
Multimodal Large Language Model, Model Compression, Model Efficiency, Post-Training Quantization
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755433isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Visual content-based indexing and retrieval
1. Introduction
@@
Figure 1. (a) Prefill visual tokens counts across different MLLMs as the image splits/resolution increases. (b) The activation values of visual tokens range from 20-20 to $10$, while textual tokens concentrate near 0 (rarely exceeding $\pm 0.5$.
Recent advances in large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b; Dubey et al., 2024) have led to remarkable performance on a wide range of natural language processing tasks. However, these models often struggle when dealing with non-textual data such as images or videos. Multimodal large language models (MLLMs) (Reid et al., 2024; Achiam et al., 2023a; Wang et al., 2023) address this limitation by integrating visual and textual modalities, thereby enabling comprehensive understanding and reasoning. Despite these benefits, their large parameter sizes, coupled with substantial computational demands, poses a major challenge for real-world deployment, particularly in resource-constrained or privacy-sensitive environments.
Challenges in MLLMs Quantization. Quantization has proven an effective strategy for reducing memory usage and inference costs in LLMs (Yuan et al., 2023a; Shao et al., 2023; Zhou et al., 2025), by converting high-precision parameters (e.g., FP32) into lower-bit representations (e.g., INT8). Yet, the transition from LLMs to MLLMs brings three unique difficulties: Time-to-First-Token (TTFT) Explosion. MLLMs often generate large numbers of visual tokens (e.g., patch embeddings or region proposals) with the resolution and aspect ratio of input images or videos. As shown in Fig 1(a), in models such as Qwen2-VL (Wang et al., 2024a), the number of prefill visual tokens grows as image resolution increases (detailed figure in Appendix A.6). As image or video resolution increases, the prefill visual tokens can escalate dramatically—leading to high TTFT and negatively impacting latency-sensitive tasks. Moreover, in higher-demand scenarios such as video-based tasks and multi-image dialogues, the accumulation of visual tokens becomes even more pronounced, further exacerbating TTFT growth. Although the commonly adopted per-token dynamic quantization (Bai et al., 2023b; Wang et al., 2024a; Liu et al., 2024; Chen et al., 2024c) offers flexibility to accommodate the larger activation distribution variance across different tokens in vision modal, its requirement for online token-wise scale recalculation during inference significantly amplifies this computational overhead. Moreover, recent studies (Chen et al., 2024a; Tan et al., 2024a) have underscored the inefficiencies of per-token dynamic quantization, especially for resource-constrained edge devices, where such methods exacerbate latency and memory pressures. Disparate Modality Distributions. As shown in Fig 1 (b), the activation distributions between visual and textual tokens reveal substantial numerical discrepancies, where visual token activations can span a significantly broader range (e.g., 20-20 to $10$) than textual tokens, which typically center near 0. A single global scale factor for both visual and textual tokens leads to either aggressive clipping of visual outliers or increased quantization granularity for text, harming overall accuracy. Visual Token Outliers. High-magnitude outliers in visual tokens, often reflecting salient image regions, make traditional clipping-based methods unsuitable. As shown in Table 1, naive clipping outliers significantly degrades accuracy, highlighting the necessity for judicious outlier treatment.
Table 1. Ablations on the clip range for different tokens.
@@
To this end, we propose MQuant, an accurate and efficient post-training quantization (PTQ) solution specifically designed for multimodal large language models (MLLMs). First, to reduce the TTFT while maintaining accuracy, we introduce Modality-Specific Static Quantization (MSQ), which apply distinct static per-tensor scales for visual vs. textual tokens based on the distinct distribution differences. Second, for further acceleration, we design an Attention-Invariant Flexible Switching (AIFS) scheme. AIFS transforms mixed multimodal tokens into unified, modality-decoupled tokens without the need for dynamic position vectors, maintaining computational equivalence and memory efficiency by avoiding the overhead associated with dynamic processing. Based on MSQ and AIFS, MQuant slashes TTFT and preserves accuracy. Third, we reveal the weight outliers caused by the online Hadamard rotations through theoretical analysis and propose Rotation Magnitude Suppression (RMS) to mitigate them. Our theoretical analysis reveals the emergence of large mean channels in MLLM weights, and RMS effectively reduces these outliers with minimal overhead.
We evaluate MQuant on five mainstream MLLMs, including InternVL (Chen et al., 2024b), Qwen-VL (Bai et al., 2023b), MiniCPM-V (Yao et al., 2024), CogVLM2 (Hong et al., 2024) and Qwen2-VL (Wang et al., 2024a). The extensive results on diverse multimodal reasoning tasks demonstrate that MQuant achieves less than 1% accuracy loss on all MLLMs with 23% and 100% speedup in prefill and decode stage under the W4A8 setting, highlighting its superior performance and wide practical value. Our main contributions are summarized as follows:
  • Insightful Observation. To our best knowledge, we present the first comprehensive analysis of quantization issues in MLLMs, revealing the root causes of performance collapse and identifying inference speed bottlenecks as well as the quantization challenges posed by modality differences.
  • MSQ and AIFS. We design Modality-specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS) to accelerate inference for heterogeneous variable-sequence multimodal inputs while maintaining accuracy.
  • RMS. We identify weight outliers caused by online rotation, and propose Rotation Magnitude Suppression to enhance quantization performance.
  • MQuant. We propose MQuant, a general PTQ framework designed for MLLMs, demonstrating both near-lossless performance and with significant speedup.
2. Preliminaries and Related Work
2.1. Multimodal Large Language Models
Figure 2. MLLM’s architecture.
As shown in the Fig. 2, the existing MLLMs framework  (Bai et al., 2023b; Chen et al., 2024c) mainly consists of three modules: (1) Visual encoder $\bm{E}$ for processing visual inputs, (2) Vision-language projector $\bm{P}$ for aligning the text and visual modalities, (3) Large language model $LLM$ that handles the multi-modal tokens and performs reasoning.
@@ 2.1. Multimodal Large Language Models
where $l$ denotes the length of $\bm{O}_{a}$. The parameter sizes of large language models (LLMs) range from 3 billion to tens of billions. Commonly used open-source LLMs include the Llama series (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a), InternLM (Cai et al., 2024), MiniCPM (Hu et al., 2024b), ChatGLM (GLM et al., 2024).
@@
MLLMs’ foundational components have greatly benefited from rapid advancements in LLM technology. Flamingo (Alayrac et al., 2022) pioneered connecting pre-trained visual encoders to LLMs, demonstrating strong generalization across visual-language tasks. Following ChatGPT (Achiam et al., 2023b), numerous open-source models based on pre-trained LLMs, such as LLaMA series (Touvron et al., 2023a, b; Dubey et al., 2024), have been proposed (Li et al., 2023a; Huang et al., 2024; Zhu et al., 2023; Liu et al., 2024). Subsequent efforts, including Qwen-VL (Bai et al., 2023b), InternVL (Chen et al., 2024c), and CogVLMV2 (Hong et al., 2024), have enhanced MLLMs by improving high-resolution inputs and scaling training data. Besides, smaller MLLMs like Mini-Gemini (Li et al., 2024c), MobileVLM (Chu et al., 2024), and MiniCPM-V (Yao et al., 2024) have emerged. Despite those advancements, the large parameter sizes of MLLMs lead to high computational costs, yet dedicated quantization methods to reduce memory usage and accelerate inference still remain underexplored.
2.2. PTQ for LLMs/MLLMs
@@
PTQ (Zhou et al., 2024; Jiang et al., 2024; Yu et al., 2025) serves as a potential strategy for model compression. By converting the high-precision variables of pre-trained models into low-bit integers, it achieves memory reduction and inference speed acceleration. For uniform quantization, given a floating-point (FP) tensor $x$ (weights or activations), it can be uniformly quantized to $b$-bits in signed quantization as follows:
(2) $\hat{\mathbf{x}}=\mathbf{Q}_{U}(\mathbf{x},b)=(clamp(\lfloor\frac{\mathbf{x}}{s}\rceil+z,q_{min},q_{max})-z)\cdot s$
where $s=\frac{\max(|\mathbf{x}|)}{2^{b-1}-1}$ is the scale factor, $\lfloor\cdot\rceil$ is the rounding-to-nearest operator, and the function $clamp(\cdot)$ clips values outside the integer range $\left[q_{min},q_{max}\right]$. $z$ is zero-point. $s$ denotes the quantization scale factor, which reflects the proportional relationship between FP values and integers. $\left[q_{min},q_{max}\right]$ is the quantization range determined by the bit-width $b$. Generally, when we quantize the network’s weight with 4-bit and activations with 8-bit, called it as W4A8. We can calculate $s$ offline using the activations from calibration samples, known as static quantization. We can also use the runtime statistics of activations to get $s$, referred to as dynamic quantization. More details are in Appendix A. 12.
Existing post-training quantization (PTQ) methods for LLMs are categorized into weight-only and weight-activation quantization (Yuan et al., 2024; Li et al., 2023b, 2024b; Yue et al., 2024; Hu et al., 2024a, 2025; Chen et al., [n. d.]; Xu et al., 2025, [n. d.]). Weight-only methods like GPTQ (Frantar et al., 2022), QuIP (Chee et al., 2024), and AWQ (Lin et al., 2023) achieve high compression rates but offer limited inference acceleration. In contrast, weight-activation quantization methods (Xiao et al., 2022; Wei et al., 2022; Yuan et al., 2023b; Zhang et al., 2024) quantize both weights and activations, improving memory usage and latency. The main challenge is activation outliers causing quantization errors. Techniques like SmoothQuant (Xiao et al., 2022) shift quantization difficulty from activations to weights, while OmniQuant (Shao et al., 2023) optimizes performance by training quantization parameters. SliceGPT (Ashkboos et al., 2024a) reduces memory demands by designing a Pre-LN + Rotate Scheme for LLMs sparsification based on computational invariance. Recently, Quarot (Ashkboos et al., 2024b) introduced applying offline Hadamard transforms and a partially online Hadamard transform to eliminate outliers, achieving state-of-the-art quantization results on LLMs. However, this solution is not applicable to MLLMs due to inherent modality differences. Building on the above studies, research on PTQ specifically tailored for MLLMs remains notably scarce. Existing studies, such as Q-VLM (Wang et al., 2024b), propose a block-wise quantization framework that optimizes cross-layer dependencies via entropy-guided partitioning. MBQ (Li et al., 2024a) uses gradients of the supervised fine-tuning loss with respect to vision and language tokens as sensitivity indicators to balance reconstruction loss during calibration. Despite efforts, existing methods still employ per-token dynamic quantization for activation, which inherently introduces a significant computational overhead than per-tensor static quantization discussed earlier. In contrast, our MQuant framework addresses these challenges by extending full static quantization to both weights and activations, specifically designed to unleash the inference potential of MLLMs while overcoming the modality-specific limitations faced by LLM-focused PTQ methods, which are not directly applicable to MLLMs due to significant modality differences.
3. Method
In this section, we present MQuant, a post-training quantization solution specifically designed for MLLMs. In Sec. 3.1, we describe proposed modality-specific static quantization (MSQ) and attention-invariant flexible switching (AIFS) mechanisms. In Sec. 3.2, we identify the weight outliers caused by the online Hadamard rotations and state Rotation Magnitude Suppression (RMS). We provide the detailed MQuant algorithm for FP MLLMs in Appendix A.1 Algorithm 1.
3.1. Modality-Specific Static Quantization and Attention-Invariant Flexible Switching
We address this issue with a two-part method: (1) Modality-Specific Static Quantization (MSQ) and (2) Attention-Invariant Flexible Switching (AIFS). MSQ applies different static scaling factors for visual and textual tokens, while AIFS reorders their positions to avoid irregular data slicing. MSQ+AIFS retains accuracy while eliminating the expensive token-wise scale online computation during inference. This approach achieves efficiency gains by leveraging per-tensor static quantization, especially for high-resolution inputs.
Figure 3. Overview of Modality-Specific Static Quantization (MSQ) and Attention-Invariant Flexible Switching (AIFS). AIFS reorders tokens so that all visual tokens appear first, then the textual tokens, while adjusting the causal mask to preserve the original model logic.
Modality-Specific Static Quantization (MSQ)
(3) $E\;=\;\underbrace{(e^{t}_{1},\dots,e^{t}_{m-1})}_{\textstyle\text{text scale }s_{t}}\;\;\underbrace{(e^{v}_{m},\dots,e^{v}_{n})}_{\textstyle\text{visual scale }s_{v}}\;\;\underbrace{(e^{t}_{n+1},\dots,e^{t}_{L})}_{\textstyle\text{text scale }s_{t}}.$
The scaling factors $s_{v}$ and $s_{t}$ correspond to visual and textual tokens, respectively. Our innovation lies in performing single-shot calibration of these factors before inference, thereby eliminating per-token computational overhead. This per-tensor static quantization strategy provides three distinct advantages: (1) complete avoidance of dynamic scale updates, (2) hardware-friendly implementation benefiting from simplified quantization kernels, and (3) prevention of distribution saturation where high-magnitude visual activations could overwhelm the narrower dynamic range of textual tokens.
Attention-Invariant Flexible Switching (AIFS)
Despite the benefits of MSQ, it complicates data handling if textual and visual tokens remain interleaved in $E$. Naïve slicing and concatenating at runtime can increase memory traffic and reduce efficiency for GEMM-based layers like QK and FC. To overcome this, we propose Attention-Invariant Flexible Switching (AIFS). Figure 3 illustrates the idea: we reorder the sequence so that all visual tokens appear first, followed by textual tokens. We then alter the causal mask to preserve the original auto-regressive relationships. Causal attention (Vaswani et al., 2017) ensures each token can only attend to earlier tokens. Formally, for a naive sequence $E$, the attention matrix is
@@ 3.1. Modality-Specific Static Quantization and Attention-Invariant Flexible Switching
With AIFS mechanism, we can reorder $E$ and obtain an unified causal mask $M^{u}_{i,j}$, the new $E^{u}=\{\;e^{v}_{m},\dots,e^{v}_{n},\;e^{t}_{1},\dots,e^{t}_{m-1},\,\dots,e^{t}_{n+1},$ $\dots,e^{t}_{L}\}$, and $M^{u}_{i,j}$ could be formulated as:
[…]
The position embeddings are also shifted accordingly (see Appendix A.2), preserving numerical equivalence of attention scores with the original sequence. This maintains auto-regressive consistency while enabling streamlined memory access: all visual/textual tokens can now be processed through respective static-scale multiplications.
Seamless Integration with FlashAttention
MSQ and AIFS, which optimizes MLLMs through optimizing on input token efficiency, while FlashAttention (Dao et al., 2022) accelerates attention computation via reduced memory access and tiling. These approaches are conceptually orthogonal, suggesting seamless integration with additive benefits without inherent conflicts. Specifically, at the beginning of inference, MSQ+AIFS reorders the intermixed multimodal tokens and passes the original visual modality start/end indices ($m$ and $n$) to FlashAttention to ensure that irrelevant positions are correctly masked. Detailed results are in Table  9.
Efficiency and Practical Benefits
Our experiments in Section 4 show that MSQ + AIFS delivers three main advantages: High Compatibility and Strong Theoretical Equivalence: The causal-mask transformation ensures that the model’s output is numerically the same as if tokens were not reordered. This reordering can be integrated into existing MLLM implementations without major code changes. Reduced Inference Latency: We replace per-token dynamic quantization with a static per-modality approach. This removes the need to recalculate scales per token, cutting runtime overhead and boosting throughput. Enhanced Memory Efficiency: Sequential processing of visual (via $s_{v}$) and textual ($s_{t}$) tokens with dedicated scale factors minimizes memory fragmentation, yielding up to 24.7% speedups and 152.9% memory efficiency gains (Table 17) through eliminated padding/slicing operations.
@@
In summary, MSQ + AIFS offers a straightforward and effective way to handle the unique challenges of multi-modal token distributions. It is valuable on edge devices or other constrained platforms where dynamic quantization is hard to implement at scale, yet per-tensor static quantization is hardware-friendly and fast (Tan et al., 2024b). More discussions is in Appendix A.10
3.2. FHT-induced Weight Outlier Mitigation: Rotation Magnitude Suppression
[…]
(6) $|W_{ij}|\;=\;\bigl{|}\,e_{i}^{\top}W\,e_{j}\bigr{|}\;\;\leq\;\;\mu\,\frac{\|W\|_{F}}{\sqrt{mn}},$